, Let k be an arbitrary natural number and p a prime number if k2 is a multiple of p then necessarily;

A.     K is an odd number

B.     K is a prime number

C.     K is a multiple of P

D.     None of the above


1
Expert's answer
2021-09-16T00:49:25-0400

If "p|k^2," then we have that.

"V_p(k^2) = 2m, where\\ 
m \in \Z^+\\"

Taking the square root of K, then

"V_p(k)=\frac{V_p(k^2)}{2}=m \in \Z^+"


Therefore, k is a multiple of p